OpenStudy (anonymous):

What are the next two numbers in this sequence? 12,14,18,26,42, ? , ? A. 64,128 B. 64,138 C. 74,128 D. 74,138

OpenStudy (jackthegreatest):

d

OpenStudy (jackthegreatest):

the difference between the numbers r 2,4,8,16,32,64...

OpenStudy (anonymous):

difference is 2^n

OpenStudy (anonymous):

Oh ok I see it now. Thank you

OpenStudy (jackthegreatest):

welcome @KatelynnGibson1099

OpenStudy (idku):

\(\large\color{black}{ \displaystyle 12\color{red}{+2}=14 }\) \(\large\color{black}{ \displaystyle 14\color{red}{+4}=18 }\) \(\large\color{black}{ \displaystyle 18\color{red}{+8}=26 }\) \(\large\color{black}{ \displaystyle 26\color{red}{+16}=42 }\) and so on...